Earlier this afternoon, while locking PRMI, I saw a big peak at 1883.48 Hz. This comes closest to the PRM's 627.75 Hz *3, so I infer that it is the 3rd order harmonic of the PRM violin mode.
While putting this in, I noticed that my addition of ETM filters the other day (elog 10746) had gotten deleted. Koji pointed out that Foton can do this - it allows you to create and save filters that are higher than 20th order, but secretly it deletes them. I went into the filter archive and recovered the old ETM filters, and split things up. I have now totally reorganized the filters, and I have made every single optic (ETMs, ITMs, PRM, SRM, BS, MC2) all the same.
FM1 is BS 1st and 2nd harmonics, and FM6 directly below that is a generic 3rd order notch that is wide enough that it encompases 3*BS.
FM2 is the PRM 1st and 2nd order, and FM7 below it is the PRM 3rd order.
FM3 is the SRM 1st order, FM4 is the ETMs' 1st order, and FM5 is the MC2 1st and 2nd order filters.
All of these filters are triggered on if any degree of freedom is triggered. They all have a ramp time of 3 sec. We may want to consider having separate trigger options for each optic, so that we're not including the PRM notch on the ETMs, for example, and vice versa.
When all of these filters are on, according to Foton we lose 5.6 degrees of phase at 100 Hz.